Can you please help me solve using the substitution method?

x + 7y = 16
-3x + 6y = 33

Answer by ankor@dixie-net.com(22740)

In the 1st equation we can see that x has a coefficient of 1 so let's arrange
the equation, so we substitute for x:
;
x + 7y = 16
x = (16 - 7y)
:
-3x + 6y = 33
:
Substitute (16-7y) for x in the 2nd equation:
-3(16-7y) + 6y = 33
:
Multiply what is inside the brackets by -3:
-48 + 21y + 6y = 33
:
Add 48 to both sides:
21y + 6y = 33 + 48
27y = 81
:
Divide both sides by 27:
y = 81/27
:
y = +3
:
:
Find x, substitute 3 for y in the 1st equation:
x + 7(3) = 16
x + 21 = 16
:
Subtract 21 from both sides:
x = 16 - 21
x = -5
:
:
Check solutions in 2nd equation:
-3(-5) + 6(3) = 33
+15 + 18 = 33 proves our solutions are correct
:
Can you do this now?